OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters? 450 m2 486 m2 504 m2 612 m2

jimthompson5910 (jim_thompson5910):

The area of the base is (9*12)/2 = 108/2 = 54 square meters So double this to get 54*2 = 108 square meters since there are 2 bases (top and bottom base)

jimthompson5910 (jim_thompson5910):

The lateral faces are rectangles. There are 3 lateral faces The first rectangle has a base of 15 with the height of 11. So the area of this rectangle is 15*11 = 165 square meters The second rectangle has a base of 12 with the height of 11. So the area of this rectangle is 12*11 = 132 square meters The third rectangle has a base of 9 with the height of 11. So the area of this rectangle is 9*11 = 99 square meters

jimthompson5910 (jim_thompson5910):

So the lateral surface area is 165+132+99 = 396 square meters

jimthompson5910 (jim_thompson5910):

Do you know what to do from here?

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

oh nvm i understand now! thanks!!

jimthompson5910 (jim_thompson5910):

what did you get?

jimthompson5910 (jim_thompson5910):

You should have added the total base area to the lateral surface area to get the total surface area

OpenStudy (anonymous):

i mean 504

jimthompson5910 (jim_thompson5910):

108+396 = 504, so you got it